how much CuSo4x5H20 and how much H20 do you need to make 500g 10% CuSo4


1
Expert's answer
2021-02-23T04:26:15-0500

The mass of pure CuSO4 needed for the solution equals "500g\times0.1=50g"

The mass of CuSO4*5H2O can be calculated:


"50g_{CuSO_4}\times\frac{1mol_{CuSO_4}}{159.61g_{CuSO_4}}\times\frac{1mol_{CuSO_4\cdot5H_2O}}{1mol_{CuSO_4}}\times\frac{249.69g_{CuSO_4\cdot5H_2O}}{1mol_{CuSO_4\cdot5H_2O}}=78g_{CuSO_4\cdot5H_2O}"


Therefore, the mass of H2O equals "500-78=422g"


Answer: 78 g CuSO4*5H2O and 422 g H2O